Problem 3-14 Schedule of Cost of Goods Manufactured; Overhead Analysis [LO3-3, LO3-4]

Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $124,600 of manufacturing overhead for an estimated allocation base of $89,000 direct material dollars to be used in production. The company has provided the following data for the just completed year:

Purchase of raw materials $ 138,000
Direct labor cost $ 85,000
Manufacturing overhead costs:
Indirect labor $ 142,400
Property taxes $ 8,300
Depreciation of equipment $ 16,000
Maintenance $ 12,000
Insurance $ 9,500
Rent, building $ 32,000
Beginning Ending
Raw Materials $ 27,000 $ 11,000
Work in Process $ 47,000 $ 37,000
Finished Goods $ 73,000 $ 61,000

Required:

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $37,000 ending balance in Work in Process includes $8,300 of direct materials. Given this assumption, supply the information missing below:
Answer #1

Solution

Gitano Products

  1. Computation of the predetermined overhead rate for the year:

Predetermined overhead rate = estimated overhead cost/direct material cost

Estimated overhead cost = $124,600

Direct material cost = $89,000

Predetermined overhead rate = (124,600/89,000) x 100 = 140%

  1. Computation of the amount of underapplied or overapplied overhead for the year:

Applied overhead –

= 140% x direct materials used

= 140% x (beginning raw materials + purchases – ending raw materials)

= 140% x (27,000 + 138,000 – 11,000) = 140% x 154,000

= $215,600

Applied overhead cost = $215,600

Actual overhead cost incurred –

Indirect labor        $142,400

Property taxes       8,300

Depreciation of equipment 16,000

Maintenance         12,000

Insurance 9,500

Rent, building       32,000

Total                      $220,200

Since actual overhead cost, $220,200 is higher than the applied overhead cost, $215,600, the overhead is underapplied.

Under applied overhead = 220,200 – 215,)600 = $4,600

  1. Schedule of cost of goods manufactured:

Gitano Products

Schedule of cost of goods manufactured:

Cost of direct materials used

$154,000

Direct labor

$85,000

Manufacturing overhead applied

$215,600

Total manufacturing costs

$454,600

Add: beginning work in process

$47,000

$501,600

Less: ending work in process

$37,000

Cost of goods manufactured

$464,600

  1. Computation of the unadjusted cost of goods sold for the year (excluding under or over applied overhead):

Gitano Products

Cost of goods sold schedule:

Cost of goods manufactured

$464,600

Add: beginning finished goods

$73,000

Cost of goods available for sale

$537,600

Less: ending finished goods

$61,000

Cost of goods sold

$476,600